OpenStudy (anonymous):

find the conjugate of the complex number 5+2i

OpenStudy (unklerhaukus):

The complex conjugate of a complex number \(z=a+ib\) is \(z^* = a-ib\). In-other-words, just negate the complex component .

OpenStudy (anonymous):

5-2i?

OpenStudy (unklerhaukus):

Yes, that's it.

OpenStudy (anonymous):

Thanks

OpenStudy (unklerhaukus):

(not too hard eh?)

OpenStudy (anonymous):

nope easy:)

OpenStudy (unklerhaukus):

\[\ddot\smile\]